OpenStudy (superhelp101):

Solve for x. 7^(x-2)=12^(x)

OpenStudy (superhelp101):

@dumbcow any help :))

OpenStudy (dumbcow):

use logs \[\ln a^n = n \ln a\]

OpenStudy (superhelp101):

can you explain that ? :/

OpenStudy (dumbcow):

its a log property you will use to solve this whenever the exponent "x" is a variable, you can take the log and bring the exponent outside of the log

OpenStudy (superhelp101):

oh okay, makes sense

OpenStudy (superhelp101):

i am a bit stuck. Can you show me ?

OpenStudy (dumbcow):

\[\rightarrow (x-2) \ln 7 = x \ln 12\]

OpenStudy (superhelp101):

would I then divide each side by ln12 to get x ?

OpenStudy (dumbcow):

no because there is an "x" on other side too distribute, then put "x" terms on same side factor out the "x" , then you divide

OpenStudy (superhelp101):

how do i do that ?

OpenStudy (dumbcow):

just like how i described :) do you know how to distribute?

OpenStudy (superhelp101):

distribute what ?

OpenStudy (superhelp101):

(x-2) ln7 ?

OpenStudy (dumbcow):

yes ... sorry got to go

OpenStudy (superhelp101):

xln(7)-2ln(7)

OpenStudy (superhelp101):

wait but what do I do next..... :((

OpenStudy (superhelp101):

@mathmale can you plzz continue from where dumbcow left off, he had to go ?

OpenStudy (mathmale):

7^(x-2)=12^(x) became \[\rightarrow (x-2) \ln 7 = x \ln 12\] (This is correct. Any questions regarding what's happening here?)

OpenStudy (superhelp101):

no

OpenStudy (mathmale):

Our job is to solve for x. In other words, we want to isolate x on the left side of this equation and have everything else on the right side.

OpenStudy (superhelp101):

okay

OpenStudy (mathmale):

Please multiply out the left side of \[\ (x-2) \ln 7 = x \ln 12\]

OpenStudy (superhelp101):

xln(7)-21ln(7)=xln(12)

OpenStudy (mathmale):

Which of these terms contain a variable and which does/do not?

OpenStudy (mathmale):

Collect all with x in them on the left; collect all constants on the right.

OpenStudy (superhelp101):

so I need to add 21ln(7) and subtract xlm(12) from both sides

OpenStudy (mathmale):

Yes, that's right. Try it.

OpenStudy (superhelp101):

ok so: xln(7)-xln(12)=21ln(7) But what happens next here ?

OpenStudy (mathmale):

Both terms on the left have x in them. Factor out x and write the remaining factor inside parentheses. Leave the right side alone.

OpenStudy (superhelp101):

ok, that makes sense, but I do I factor the x ?

OpenStudy (superhelp101):

@mathmale

OpenStudy (mathmale):

xln(7)-xln(12)=21ln(7) becomes x (ln 7 - ln 12) = 21 ln 7.

OpenStudy (superhelp101):

oh okay

OpenStudy (mathmale):

How would you now solve for x?

OpenStudy (superhelp101):

do i use a calculator and solve the inside parenthesis first ?

OpenStudy (mathmale):

No. Best to use a certain "rule of logs" to combine ln 7 - ln 12 into one log. Of course, if you're supposed to find the answer as a decimal fraction, then go ahead and use your calculator.

OpenStudy (superhelp101):

ok how would i combine, can you teach me it ?

OpenStudy (mathmale):

The rule of logs that applies here is log a - log b = log (a/b). Therefore, log 7 -log 12 = ?

OpenStudy (superhelp101):

log(7/12)

OpenStudy (superhelp101):

ok I got it from there, thank youuuuuu...hhehe :)))) What are the benefits from getting the membership on os. Do you like it ?

OpenStudy (mathmale):

You're welcome! I'm a retired teacher and love to continue working with students on their math through OpenStudy. I'm also a Moderator and can thus help keep OS clean and focused.

OpenStudy (superhelp101):

That's awesome! :) you're very kind! :)

OpenStudy (superhelp101):

wait, but I try to solve x and I get -16.625 and when I plug it in it doesn't seem to work

OpenStudy (superhelp101):